OpenStudy (anonymous):

How do I find an equation of the line containing the centers of the following two circles? x2 +y2-4x+6y+4=0 and x2+y2+6x+4y+9=0

OpenStudy (amistre64):

find the coordinates for the centers of each circle and find the slope between them and plug them into the line equation :)

OpenStudy (amistre64):

(x^2-4x+n) + (y^2 +6y +m) = -4 +n +m (x^2 -4x +4) + (y^2 +6x +9) = -4 +4 +9 (x-2)^2 + (y+3)^2 = sqrt(9) = 3 center is at (x=2, y=-3) (2,-3)

OpenStudy (amistre64):

(x^2 +6x +n)+(y^2 +4y +m) = -9+n+m (x^2 +6x +9) + (y^2 +4x +4) = -9 +9 +4 (x+3)^2 + (y+2)^2 = sqrt(4) = 2 center of circle is at (x=-3, y=-2) (-3,-2)

OpenStudy (amistre64):

(2,-3) (-3,-2) m = (y1 -y2)/(x1-x2) m= -3--2/2--3 = -3+2/2+3 = -1/5 y = mx+b y = (-1/5)x + b -2=(-1/5)(-3) + b -2 = 3/5 + b -2 - 3/5 = b -13/5 = b y = (-1/5)x - 13/5 OR x + 5y = 13

OpenStudy (amistre64):

might want another look at that....

OpenStudy (amistre64):

x + 5y = -13 is better :)